OpenStudy (anonymous):

Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 10 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. Part B: How many cubic inches of water was poured into the container? Part C: What was the height of the water in the container? Show your work. Measurement is given as length by width by height.

OpenStudy (alekos):

first convert the 10cm to inches then work out the max volume for part 1

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

will help give me 10 mins taking a test

OpenStudy (thehumantorch10):

Part A: V = L * W * H V = 10" * 6" * 9 cm (convert 9cm to inches = 9 * 2.54 = 22.86) V = 10 * 6 * 22.86 V = 1371.6 Now all measurement are in the same unit. Part B: 1 cup is = 14.44 cubic inches So, 14.44 * 15 = 216.6 216.6 cubic inches of water have been poured into the container. Part C: The maximum volume of water that the rectangular container can hold is V = 1371.6 Mike poured 216.6 cubic inches into the pan. 216.6 = 10 * 6 * H 216.6 / 60 = 3.61

OpenStudy (anonymous):

thnx :)